OpenStudy (anonymous):

The force needed to keep a car from skidding on a curve varies inversely with the radius of the curve and jointly with the weight of the car and the square of the speed. If 4500 pounds of force keeps a 2500 pound car from skidding as it travels at 45 miles per hour on a curve of radius 1000 feet, what force would keep the car from skidding on a curve of radius 800 feet at 60 miles per hour?

OpenStudy (kropot72):

\[4500=k \times \frac{2500\times 45^{2}}{1000}\] where k is a constant of proportionality. \[k=\frac{4500\times 1000}{2500\times 45^{2}}=0.88888888\] The required force is given by: \[F=\frac{0.8888888\times 2500\times 60^{2}}{800}=you\ can\ calculate\]